simplify complex fractions
(-a+b/3ab) / (a-b/2ab)
many thanks

Answer by dabanfield(803)

(-a+b/3ab) / (a-b/2ab)
Invert and multiply:
((-a+b)/3ab))*(2ab/(a-b))=
((-a+b)*2ab))/(3ab*(a-b))=
(-(a-b)/(a-b))*(2ab/3ab) =
-1*2/3 = -2/3